0 0
Add a comment Improve this question Transcribed image text
Answer #1

apply in atransformer N1/N2 = V1/V2 = i2/i1


so N1/N2 = 231/121 = 1.909

---------------------------------

i2 = 14* 121/231

i2 = 7.33 Amps